a company manufactures two kinds of ice-cream. the vanilla ice-cream sells for $2.50 each while the chocolate flavour ice-cream seels for $4.50 cents each. it costs the company 1 labour hour to make the vanilla ice-cream and 2 labour hours to make the chocolate flavour ice-cream. the company has a total of 300 labour hours available. it costs the company 3 machine hours for the vanilla ice-cream and 2 machine hours for the chocolate flavour ice-cream. the company has a total of 480 machine hours available. how much of each type of ice-cream should the company produce to maximize revenue? what is the maximum revenue? [ hint: let vanilla ice-cream = x]
1
Expert's answer
2020-05-08T19:41:32-0400

Let vanilla ice creams produced be x and chocolate ice cream be represented by y.

So,Max Z=2.50x+4.50y

subject to:

x+2y "\leq" 300 (Labour hour constraint)

"3x+2y\leq 400" (Machine hour constraint)

x"\geq0, y\geq 0"

Now,plotting it on graph:

Labour hour constraint is red line and Machine hour constraint is blue line.



Now finding values at the feasible points.

Points Value

(0,0) 0

(160,0) 400

(90,105) 697.5

(0,150) 675


So,90 vanilla and 105 chocolate ice cream to be made for obtaining maximum revenue which is $697.5